# Two-dimensional FrBD friction models for rolling contact

### v1

@@ 2.1.1 Rheological model
First, the situation depicted in Fig. 1(a) is considered: the upper body travels with rigid relative velocity $\mathbb{R}^{2}\ni\bm{v}_{\textnormal{r}}=[v_{\textnormal{r}x}\;v_{\textnormal{r}y}]^{\mathrm{T}}$ with respect to the rigid substrate (lower body). It should be noted that, in Fig. 1, the rigid relative velocity $\bm{v}_{\textnormal{r}}$ is defined positive with a minus sign in the reference frame $(O;x,y,z)$; this simplifies the notation for the final formulae and ensures notational consistency with other dynamic friction models such as the Dahl and LuGre [58, 59, 60]. To the lower boundary of the upper body, deformable massless bristles protrude, whose deflection is denoted by $\mathbb{R}^{2}\ni\bm{z}=[z_{x}\;z_{y}]^{\mathrm{T}}$ (not to be confounded with the vertical axis $z$). The roots of the bristles, which are attached to the upper body, have rigid relative velocity $\bm{v}_{\textnormal{r}}$ with respect to the substrate. On the other hand, the total sliding velocity between the tip of the bristle and the lower body reads
$\bm{v}_{\textnormal{s}}(\dot{\bm{z}},\bm{v}_{\textnormal{r}})=\bm{v}_{\textnormal{r}}-\dot{\bm{z}},$ (1)
and is again defined as positive with a minus sign convention.
Figur 1: A schematic representation of the friction model: (a) configuration with a rigid substrate; (b) configuration with a deformable substrate. The problem is studied in a reference frame $(O;x,y,z)$ with unit vectors $(\hat{\bm{e}}_{x},\hat{\bm{e}}_{y},\hat{\bm{e}}_{z})$.
@@ 2.1.2 Local friction model and friction dissipation
$\begin{split}\bm{f}_{\textnormal{r}}^{\star}&=\operatorname{argmax}_{\bm{f}_{\textnormal{r}}\in\mathcal{C}(\bm{v}_{\textnormal{s}}(t))}\bm{f}_{\textnormal{r}}^{\mathrm{T}}\bm{v}_{\textnormal{s}}(t),\qquad\bm{v}_{\textnormal{s}}(t)\in\mathbb{R}^{2},\;t\in[0,T],\end{split}$ (10)
@@
$\bm{f}_{\textnormal{r}}^{\star}\bigl(\bm{v}_{\textnormal{s}}(t)\bigr)=\dfrac{\mathbf{M}^{2}\bigl(\bm{v}_{\textnormal{s}}(t)\bigr)\bm{v}_{\textnormal{s}}(t)}{\mathinner{\!\left\lVert\mathbf{M}\bigl(\bm{v}_{\textnormal{s}}(t)\bigr)\bm{v}_{\textnormal{s}}(t)\right\rVert}_{2}},\qquad\bm{v}_{\textnormal{s}}(t)\in\mathbb{R}^{2}\setminus\{\bm{0}\},\;t\in[0,T],$ (14)
@@
$\bm{f}_{\textnormal{r}}^{\star}\bigl(\bm{v}_{\textnormal{s}}(t)\bigr)^{\mathrm{T}}\bm{v}_{\textnormal{s}}(t)=\mathinner{\!\left\lVert\mathbf{M}\bigl(\bm{v}_{\textnormal{s}}(t)\bigr)\bm{v}_{\textnormal{s}}(t)\right\rVert}_{2},\qquad t\in[0,T].$ (15)
@@
$\bm{f}_{\textnormal{r}}^{\star}(\bm{v}_{\textnormal{s}})=\dfrac{\mathbf{M}^{2}(\bm{v}_{\textnormal{s}})\bm{v}_{\textnormal{s}}}{\mathinner{\!\left\lVert\mathbf{M}(\bm{v}_{\textnormal{s}})\bm{v}_{\textnormal{s}}\right\rVert}_{2,\varepsilon}},$ (16)
@@
$\bm{f}_{\textnormal{r}}^{\star}(\bm{v}_{\textnormal{s}})=\mu(\bm{v}_{\textnormal{s}})\dfrac{\bm{v}_{\textnormal{s}}}{\mathinner{\!\left\lVert\bm{v}_{\textnormal{s}}\right\rVert}_{2}},\qquad\bm{v}_{\textnormal{s}}\in\mathbb{R}^{2}\setminus\{\bm{0}\}.$ (17)
@@
$\bm{f}_{\textnormal{r}}^{\star}(\bm{v}_{\textnormal{s}})=\mu(\bm{v}_{\textnormal{s}})\dfrac{\bm{v}_{\textnormal{s}}}{\mathinner{\!\left\lVert\bm{v}_{\textnormal{s}}\right\rVert}_{2,\varepsilon}},$ (18)
@@ 2.2 Dynamic friction model derivation
Starting with Eqs. (2) and (16), Theorem (2.2) may be sapiently invoked to generalise the FrBD model to the two-dimensional contact case. The fundamental observation is that, in the absence of inertial effects, the net bristle force $\bm{f}(\dot{\bm{z}},\bm{z})$ should be equal to the frictional one $\bm{f}_{\textnormal{r}}(\bm{v}_{\textnormal{s}}(\dot{\bm{z}},\bm{v}_{\textnormal{r}}))$ generated at the bodies’ interface. Therefore, adopting the notation of Theorem 2.2 and equating Eqs. (2) and (16) gives
@@
$\begin{split}\nabla_{\bm{\dot{z}}}\bm{H}(\dot{\bm{z}},\bm{z},\bm{v}_{\textnormal{r}})^{\mathrm{T}}&=\nabla_{\dot{\bm{z}}}\bm{f}(\dot{\bm{z}},\bm{z})^{\mathrm{T}}-\nabla_{\dot{\bm{z}}}\bm{v}_{\textnormal{s}}(\dot{\bm{z}},\bm{v}_{\textnormal{r}})^{\mathrm{T}}\nabla_{\bm{v}_{\textnormal{s}}}\bm{f}_{\textnormal{r}}\bigl(\bm{v}_{\textnormal{s}}(\dot{\bm{z}},\bm{v}_{\textnormal{r}})\bigr)^{\mathrm{T}}\\ &=\nabla_{\dot{\bm{z}}}\bm{f}(\dot{\bm{z}},\bm{z})^{\mathrm{T}}+\nabla_{\bm{v}_{\textnormal{s}}}\bm{f}_{\textnormal{r}}\bigl(\bm{v}_{\textnormal{s}}(\dot{\bm{z}},\bm{v}_{\textnormal{r}})\bigr)^{\mathrm{T}}\approx\nabla_{\dot{\bm{z}}}\bm{f}(\dot{\bm{z}},\bm{z})^{\mathrm{T}}+\dfrac{\mathbf{M}^{2}\bigl(\bm{v}_{\textnormal{s}}(\dot{\bm{z}},\bm{v}_{\textnormal{r}})\bigr)}{\mathinner{\!\left\lVert\mathbf{M}\bigl(\bm{v}_{\textnormal{s}}(\dot{\bm{z}},\bm{v}_{\textnormal{r}})\bigr)\bm{v}_{\textnormal{s}}(\dot{\bm{z}},\bm{v}_{\textnormal{r}})\right\rVert}_{2,\varepsilon}},\end{split}$ (22)
@@
$\dot{\bm{z}}(t)=-\mathbf{G}^{-1}\bigl(\bm{v}_{\textnormal{r}}(t)\bigr)\biggl[\mathbf{\Sigma}_{0}\mathinner{\!\left\lVert\mathbf{M}\bigl(\bm{v}_{\textnormal{r}}(t)\bigr)\bm{v}_{\textnormal{r}}(t)\right\rVert}_{2,\varepsilon}\bm{z}(t)-\mathbf{M}^{2}\bigl(\bm{v}_{\textnormal{r}}(t)\bigr)\bm{v}_{\textnormal{r}}(t)\biggr],\qquad t\in(0,T),$ (23a)
@@
Equations (23) and (24) describe the two-dimensional FrBD model. In the sequel, it will be particularised to describe various distributed contact phenomena, including rolling and spinning. Before moving to Sect. 3, it is beneficial to draw some preliminary considerations about the passive and dissipative nature of the FrBD model, as well as on its mathematical structure. More specifically, regarding $\bm{v}_{\textnormal{s}}(\dot{\bm{z}},\bm{v}_{\textnormal{r}})$ and $\bm{f}(\dot{\bm{z}},\bm{z})$ as the input and otuput of the model, and combining Eqs. (2) and (15) yields
$\begin{split}\bm{f}_{\textnormal{r}}^{\star}\bigl(\bm{v}_{\textnormal{s}}(\dot{\bm{z}},\bm{v}_{\textnormal{r}})\bigr)^{\mathrm{T}}\bm{v}_{\textnormal{s}}(\dot{\bm{z}},\bm{v}_{\textnormal{r}})=\bm{f}(\dot{\bm{z}},\bm{z})^{\mathrm{T}}\bm{v}_{\textnormal{s}}(\dot{\bm{z}},\bm{v}_{\textnormal{r}})=\bm{z}^{\mathrm{T}}\mathbf{\Sigma}_{0}\bm{v}_{\textnormal{s}}(\dot{\bm{z}},\bm{v}_{\textnormal{r}})+\dot{\bm{z}}^{\mathrm{T}}\mathbf{\Sigma}_{1}\bm{v}_{\textnormal{s}}(\dot{\bm{z}},\bm{v}_{\textnormal{r}})\geq 0.\end{split}$ (25)
@@
$\bm{f}(\dot{\bm{z}},\bm{z})^{\mathrm{T}}\bm{v}_{\textnormal{r}}=\bm{z}^{\mathrm{T}}\mathbf{\Sigma}_{0}\bm{v}_{\textnormal{r}}+\dot{\bm{z}}^{\mathrm{T}}\mathbf{\Sigma}_{1}\bm{v}_{\textnormal{r}}\geq\bm{z}^{\mathrm{T}}\mathbf{\Sigma}_{0}\dot{\bm{z}}+\dot{\bm{z}}^{\mathrm{T}}\mathbf{\Sigma}_{1}\dot{\bm{z}}.$ (26)
@@ 3.1.1 Eulerian approach and transport velocity
$\mathbf{H}(\bar{\bm{v}}_{\textnormal{r}},s)\triangleq\mathbf{G}^{-1}\bigl(V_{\textnormal{r}}(s)\bar{\bm{v}}_{\textnormal{r}}\bigr)\mathbf{M}^{2}\bigl(V_{\textnormal{r}}(s)\bar{\bm{v}}_{\textnormal{r}}\bigr).$ (30)
Equation (28a) is now a PDE, and more specific a transport equation. In addition to the initial condition (IC) (28b), it needs to be supplemented with a BC, as discussed next in Sect. 3.1.2.
@@ 3.2 Standard linear rolling contact model
$\bar{\bm{v}}(\bm{x},s)=\bm{\sigma}(s)+\mathbf{A}_{\varphi}(s)\bm{x},$ (39)
@@ 3.3 Semilinear rolling contact model for large spin slips
Large spin slips give rise to geometric effects that are not accurately captured by Model 3.2. The classes of models introduced below describe the dynamics of deformable bodies rolling on a rigid, flat substrate in the presence of substantial spin slip. Their primary applications include tyre–road interaction and spherical robots. The same formulations also apply to other elastic rolling contact pairs, such as wheel-rail systems, but these typically experience only modest spin slip, making the full theory developed here less relevant in practice.
@@
$\bar{\bm{v}}_{\textnormal{r}}\bigl(\bm{z}(\bm{x},s),\bm{x},s\bigr)=\bar{\bm{v}}(\bm{x},s)+\mathbf{A}_{\varphi_{2}}(s)\bm{z}(\bm{x},s),$ (48)
where $\bar{\bm{v}}(\bm{x},s)$ reads again as in (39), and $\mathbf{A}_{\varphi_{2}}(s)\in\mathbf{M}_{2}(\mathbb{R})$ reads

### v2

@@ 2.1.1 Rheological model
First, the situation depicted in Fig. 1(a) is considered: the upper body travels with rigid relative velocity $\mathbb{R}^{2}\ni\bm{v}_{\textnormal{r}}=[v_{\textnormal{r}x}\;v_{\textnormal{r}y}]^{\mathrm{T}}$ with respect to the rigid substrate (lower body). To the lower boundary of the upper body, deformable massless bristles protrude, whose deflection is denoted by $\mathbb{R}^{2}\ni\bm{z}=[z_{x}\;z_{y}]^{\mathrm{T}}$ (not to be confounded with the vertical axis $z$). These bristles, which are interpreted as material particles in the paper, consistently with the classic theories of contact mechanics [5], are assumed to have zero vertical dimension, and only undergo deformations along the $x$ and $y$ directions. The roots of the bristles, which are attached to the upper body, have rigid relative velocity $\bm{v}_{\textnormal{r}}$ with respect to the substrate. On the other hand, the total sliding velocity between the tip of the bristle and the lower body reads
$\bm{v}_{\textnormal{s}}(\dot{\bm{z}},\bm{v}_{\textnormal{r}})=\bm{v}_{\textnormal{r}}+\dot{\bm{z}}.$ (1)
Figur 1: A schematic representation of the friction model: (a) configuration with a rigid substrate; (b) configuration with a deformable substrate. The problem is studied in a right-handed reference frame $(O;x,y,z)$ with unit vectors $(\hat{\bm{e}}_{x},\hat{\bm{e}}_{y},\hat{\bm{e}}_{z})$.
@@ 2.1.2 Local friction model and friction dissipation
$\begin{split}\bm{f}_{\textnormal{r}}^{\star}&=\operatorname{argmax}_{\bm{f}_{\textnormal{r}}\in\mathcal{C}(\bm{v}_{\textnormal{s}}(t))}-\bm{f}_{\textnormal{r}}^{\mathrm{T}}\bm{v}_{\textnormal{s}}(t),\qquad\bm{v}_{\textnormal{s}}(t)\in\mathbb{R}^{2},\;t\in[0,T],\end{split}$ (10)
@@
$\bm{f}_{\textnormal{r}}^{\star}\bigl(\bm{v}_{\textnormal{s}}(t)\bigr)=-\dfrac{\mathbf{M}^{2}\bigl(\bm{v}_{\textnormal{s}}(t)\bigr)\bm{v}_{\textnormal{s}}(t)}{\mathinner{\!\left\lVert\mathbf{M}\bigl(\bm{v}_{\textnormal{s}}(t)\bigr)\bm{v}_{\textnormal{s}}(t)\right\rVert}_{2}},\qquad\bm{v}_{\textnormal{s}}(t)\in\mathbb{R}^{2}\setminus\{\bm{0}\},\;t\in[0,T],$ (14)
@@
$-\bm{f}_{\textnormal{r}}^{\star}\bigl(\bm{v}_{\textnormal{s}}(t)\bigr)^{\mathrm{T}}\bm{v}_{\textnormal{s}}(t)=\mathinner{\!\left\lVert\mathbf{M}\bigl(\bm{v}_{\textnormal{s}}(t)\bigr)\bm{v}_{\textnormal{s}}(t)\right\rVert}_{2},\qquad t\in[0,T].$ (15)
@@
$\bm{f}_{\textnormal{r}}^{\star}(\bm{v}_{\textnormal{s}})=-\dfrac{\mathbf{M}^{2}(\bm{v}_{\textnormal{s}})\bm{v}_{\textnormal{s}}}{\mathinner{\!\left\lVert\mathbf{M}(\bm{v}_{\textnormal{s}})\bm{v}_{\textnormal{s}}\right\rVert}_{2,\varepsilon}},$ (16)
@@
$\bm{f}_{\textnormal{r}}^{\star}(\bm{v}_{\textnormal{s}})=-\mu(\bm{v}_{\textnormal{s}})\dfrac{\bm{v}_{\textnormal{s}}}{\mathinner{\!\left\lVert\bm{v}_{\textnormal{s}}\right\rVert}_{2}},\qquad\bm{v}_{\textnormal{s}}\in\mathbb{R}^{2}\setminus\{\bm{0}\}.$ (17)
@@
$\bm{f}_{\textnormal{r}}^{\star}(\bm{v}_{\textnormal{s}})=-\mu(\bm{v}_{\textnormal{s}})\dfrac{\bm{v}_{\textnormal{s}}}{\mathinner{\!\left\lVert\bm{v}_{\textnormal{s}}\right\rVert}_{2,\varepsilon}},$ (18)
@@ 2.2 Dynamic friction model derivation
Starting with Eqs. (2) and (16), Theorem (2.2) may be sapiently invoked to generalise the FrBD model to the two-dimensional contact case. The fundamental observation is that, in the absence of inertial effects, the net bristle force $\bm{f}(\dot{\bm{z}},\bm{z})$ should be equal to the frictional one $\bm{f}_{\textnormal{r}}(\bm{v}_{\textnormal{s}}(\dot{\bm{z}},\bm{v}_{\textnormal{r}}))$ generated at the bodies’ interface, that is, $\bm{f}(\dot{\bm{z}},\bm{z})=\bm{f}_{\textnormal{r}}(\bm{v}_{\textnormal{s}}(\dot{\bm{z}},\bm{v}_{\textnormal{r}}))$. Therefore, adopting the notation of Theorem 2.2 and equating Eqs. (2) and (16) gives
@@
$\begin{split}\nabla_{\bm{\dot{z}}}\bm{H}(\dot{\bm{z}},\bm{z},\bm{v}_{\textnormal{r}})^{\mathrm{T}}&=\nabla_{\dot{\bm{z}}}\bm{f}(\dot{\bm{z}},\bm{z})^{\mathrm{T}}-\nabla_{\dot{\bm{z}}}\bm{v}_{\textnormal{s}}(\dot{\bm{z}},\bm{v}_{\textnormal{r}})^{\mathrm{T}}\nabla_{\bm{v}_{\textnormal{s}}}\bm{f}_{\textnormal{r}}\bigl(\bm{v}_{\textnormal{s}}(\dot{\bm{z}},\bm{v}_{\textnormal{r}})\bigr)^{\mathrm{T}}\\ &=\nabla_{\dot{\bm{z}}}\bm{f}(\dot{\bm{z}},\bm{z})^{\mathrm{T}}-\nabla_{\bm{v}_{\textnormal{s}}}\bm{f}_{\textnormal{r}}\bigl(\bm{v}_{\textnormal{s}}(\dot{\bm{z}},\bm{v}_{\textnormal{r}})\bigr)^{\mathrm{T}}\approx\nabla_{\dot{\bm{z}}}\bm{f}(\dot{\bm{z}},\bm{z})^{\mathrm{T}}+\dfrac{\mathbf{M}^{2}\bigl(\bm{v}_{\textnormal{s}}(\dot{\bm{z}},\bm{v}_{\textnormal{r}})\bigr)}{\mathinner{\!\left\lVert\mathbf{M}\bigl(\bm{v}_{\textnormal{s}}(\dot{\bm{z}},\bm{v}_{\textnormal{r}})\bigr)\bm{v}_{\textnormal{s}}(\dot{\bm{z}},\bm{v}_{\textnormal{r}})\right\rVert}_{2,\varepsilon}},\end{split}$ (22)
@@
$\dot{\bm{z}}(t)=-\mathbf{G}^{-1}\bigl(\bm{v}_{\textnormal{r}}(t)\bigr)\biggl[\mathbf{\Sigma}_{0}\mathinner{\!\left\lVert\mathbf{M}\bigl(\bm{v}_{\textnormal{r}}(t)\bigr)\bm{v}_{\textnormal{r}}(t)\right\rVert}_{2,\varepsilon}\bm{z}(t)+\mathbf{M}^{2}\bigl(\bm{v}_{\textnormal{r}}(t)\bigr)\bm{v}_{\textnormal{r}}(t)\biggr],\qquad t\in(0,T),$ (23a)
@@
Equations (23) and (24) describe the two-dimensional FrBD model. In the sequel, it will be particularised to describe various distributed contact phenomena, including rolling and spinning. Before moving to Sect. 3, it is beneficial to draw some preliminary considerations about the mathematical structure of the FrBD model, as well as on its dissipative nature. Some preliminary intuition may be gained by first considering the stationary solution to Eq. (23), which reads
$\bm{z}(\bm{v}_{\textnormal{r}})=-\mathbf{\Sigma}_{0}^{-1}\dfrac{\mathbf{M}^{2}(\bm{v}_{\textnormal{r}}\bigr)\bm{v}_{\textnormal{r}}}{\mathinner{\!\left\lVert\mathbf{M}(\bm{v}_{\textnormal{r}})\bm{v}_{\textnormal{r}}\right\rVert}_{2,\varepsilon}}.$ (25)
The corresponding stationary bristle force, obtained by setting $\dot{\bm{z}}=\bm{0}$ in Eq. (2) reads
$\bm{f}(\bm{v}_{\textnormal{r}})=\bm{f}_{\textnormal{r}}(\bm{v}_{\textnormal{r}})=-\dfrac{\mathbf{M}^{2}(\bm{v}_{\textnormal{r}}\bigr)\bm{v}_{\textnormal{r}}}{\mathinner{\!\left\lVert\mathbf{M}(\bm{v}_{\textnormal{r}})\bm{v}_{\textnormal{r}}\right\rVert}_{2,\varepsilon}}.$ (26)
Equation (26) shows that, under steady-state conditions ($\dot{\bm{z}}(t)=\bm{0}$), the FrBD model can reproduce arbitrary static friction characteristics. It is worth noting that, unlike the LuGre model, a minus sign appears in Eqs. (25) and (26). This difference is consistent with the adopted sign convention and with the formulation proposed in [76]. More generally, the ODE (23) governs the evolution of the bristle dynamics in the vicinity of sliding, that is, for relatively small deformation velocities $\dot{\bm{z}}\approx\bm{0}$ and $\bm{v}_{\textnormal{s}}(\dot{\bm{z}},\bm{v}_{\textnormal{r}})\approx\bm{v}_{\textnormal{r}}$.
The second observation pertains the passivity properties of the model. More specifically, regarding $\bm{v}_{\textnormal{s}}(\dot{\bm{z}},\bm{v}_{\textnormal{r}})$ and $\bm{f}(\dot{\bm{z}},\bm{z})$ as the input and otuput, respectively, and combining Eqs. (2) and (15) yields
$\begin{split}-\bm{f}_{\textnormal{r}}^{\star}\bigl(\bm{v}_{\textnormal{s}}(\dot{\bm{z}},\bm{v}_{\textnormal{r}})\bigr)^{\mathrm{T}}\bm{v}_{\textnormal{s}}(\dot{\bm{z}},\bm{v}_{\textnormal{r}})=-\bm{f}(\dot{\bm{z}},\bm{z})^{\mathrm{T}}\bm{v}_{\textnormal{s}}(\dot{\bm{z}},\bm{v}_{\textnormal{r}})=-\bm{z}^{\mathrm{T}}\mathbf{\Sigma}_{0}\bm{v}_{\textnormal{s}}(\dot{\bm{z}},\bm{v}_{\textnormal{r}})-\dot{\bm{z}}^{\mathrm{T}}\mathbf{\Sigma}_{1}\bm{v}_{\textnormal{s}}(\dot{\bm{z}},\bm{v}_{\textnormal{r}})\geq 0.\end{split}$ (27)
@@
$-\bm{f}(\dot{\bm{z}},\bm{z})^{\mathrm{T}}\bm{v}_{\textnormal{r}}=-\bm{z}^{\mathrm{T}}\mathbf{\Sigma}_{0}\bm{v}_{\textnormal{r}}-\dot{\bm{z}}^{\mathrm{T}}\mathbf{\Sigma}_{1}\bm{v}_{\textnormal{r}}\geq\bm{z}^{\mathrm{T}}\mathbf{\Sigma}_{0}\dot{\bm{z}}+\dot{\bm{z}}^{\mathrm{T}}\mathbf{\Sigma}_{1}\dot{\bm{z}}.$ (28)
@@ 3.1.1 Eulerian approach and transport velocity
$\mathbf{H}(\bar{\bm{v}}_{\textnormal{r}},s)\triangleq-\mathbf{G}^{-1}\bigl(V_{\textnormal{r}}(s)\bar{\bm{v}}_{\textnormal{r}}\bigr)\mathbf{M}^{2}\bigl(V_{\textnormal{r}}(s)\bar{\bm{v}}_{\textnormal{r}}\bigr).$ (32)
Equation (30a) is now a PDE, and more specifically a transport equation. In addition to the initial condition (IC) (30b), it needs to be supplemented with a BC, as discussed next in Sect. 3.1.2.
@@ 3.2 Standard linear rolling contact model
$\bar{\bm{v}}(\bm{x},s)=-\bm{\sigma}(s)-\mathbf{A}_{\varphi}(s)\bm{x},$ (41)
@@ 3.3 Semilinear rolling contact model for large spin slips
Large spin slips give rise to geometric effects that are not accurately captured by Model 3.2. The classes of models introduced below describe the dynamics of deformable bodies rolling on a rigid, flat substrate in the presence of substantial spin slip. Their primary applications include tyre-road interaction and spherical robots. The same formulations also apply to other elastic rolling contact pairs, such as wheel-rail systems, but these typically experience only modest spin slip, making the full theory developed here less relevant in practice.
@@
$\bar{\bm{v}}_{\textnormal{r}}\bigl(\bm{z}(\bm{x},s),\bm{x},s\bigr)=\bar{\bm{v}}(\bm{x},s)-\mathbf{A}_{\varphi_{2}}(s)\bm{z}(\bm{x},s),$ (50)
where $\bar{\bm{v}}(\bm{x},s)$ reads again as in (41), and $\mathbf{A}_{\varphi_{2}}(s)\in\mathbf{M}_{2}(\mathbb{R})$ writes